# Transfer Learning for T-Cell Response Prediction

### v1

@@ 3.2 Negative transfer between domains
In a per-source setting, we train separate BASE-T models for every peptide source, each of which contains peptides from several MHC alleles. To make the performance of the models in the per-source setting directly comparable to the multi-domain setting, we first aggregate the predictions for the each source (performed by different models) into one data set again. Then we apply the evaluation on groups of peptide source/MHC allele combinations (see Section 2.8) exactly as in the multi-domain setting. This evaluation also ensures that only predictions coming from the same per-source model are compared to each other because AUC values are computed per peptide source/MHC allele combination. Conversely, the per-allele setting is a combination of separate BASE-T models for each MHC allele, each of which is trained on data from several peptide sources.
The corresponding validation study (Supplement S4) shows that the AUC  of the multi-domain model is reduced by 0.05 compared to the per-source models and by 0.02 compared to the per-allele models, confirming our hypothesis that there is negative transfer between peptide sources and, to a smaller degree, between MHC alleles.
@@ 3.3 Improved predictions by per-source fine-tuning
FINE-T achieves higher AUC  values than BASE-T for all ten source/MHC combinations. We also observe that the difference is stronger for MHC I peptides than for MHC II peptides, which may be due to more room for improvement as BASE-T performs worse for MHC I compared to MHC II peptides. We also studied the performance of separate per-source models without fine-tuning. The resulting AUC values (see Supplement S5) are overall in between BASE-T and FINE-T, suggesting that the combination of both ideas, per source models and fine-tuning, are responsible for the superior performance of FINE-T.
By generating many random permutations of the T-cell response labels, we also verify that FINE-T learns a signal that cannot be explained by random structure in the data (see Supplement S6).
Figure 5: Test performance on human peptides. For both MHC classes, the first two rows correspond to existing models from the literature, which are trained on other data sets. Bag-of-AA baseline and FINE-T are trained on the same data. Blue bars correspond to the evaluation on the previously completely unused test set. Grey bars show the mean AUC from the final nested cross validation.
@@ 3.4 Comparison to other tools for human peptides
T-cell response predictions for peptides from human proteins are of special interest for the development of personalized cancer vaccines since neo-antigens also originate from human proteins (Lang et al., 2022). Thus, we focus our analysis on human peptides and compare the performance of FINE-T, which was the best model in the validation study, to existing pre-trained models from the literature.
NetMHCpan 4.1 and NetMHCIIpan 4.0 (Reynisson et al., 2020) have been trained to predict the presentation of peptides on MHC molecules. Despite not being trained on T-cell response data, we include these models since they are commonly used in the context of personalized cancer vaccines. PRIME 2.0 (Gfeller et al., 2023) has been trained on T-cell response data with MHC I neo-antigens, whereas CD4Episcore (Dhanda et al., 2018) has been trained on MHC II T-cell response data from several peptide sources. Since the training data for all pre-trained models deviates from our training data, we also include the Bag-Of-AA baseline, trained on the same data set as FINE-T, in the comparison. In contrast to all previous study, we now train a model on the union of the first four data subsets, and evaluate on the previously unused data subset $\mathcal{D}_{5}$. In addition, we also carry out a full nested cross validation. The results are shown in Figure 5.
For MHC I peptides, FINE-T achieves with a value of 0.615 the highest AUC among all models. PRIME 2.0 yields the lowest performance, which is likely due to the differences in the training data. For MHC II peptides, FINE-T also outperforms the existing approaches, and the AUC is, with a value 0.68 even slightly better than in the validation study. While Paul et al. (2015) have observed that combining MHC binding predictions from several alleles has a predictive value for immunogenicity, on the data set we use this applies only to MHC I (NetMHCpan 4.1) but not to MHC II (NetMHCIIpan 4.0). The large performance difference between the Bag-Of-AA model and FINE-T suggests that the possibility to use positional information of amino acids in the peptide sequence is indeed beneficial, and more so for MHC class II compared to MHC class I.
@@ 4 Discussion
Predicting T-cell responses to peptides is a challenging problem not only due to the limited sample size, but also since the data features a multi-domain structure, which entails, as illustrated in this article, the danger of shortcut learning.
@@
To ensure that performance estimates are not inflated by the use of shortcuts, we have defined a domain-aware evaluation scheme. We have then demonstrated that shortcut learning can be effectively reduced by applying adversarial domain adaptation. Shortcut reduction did, however, not directly improve predictive performance, which is likely due to large domain differences and negative transfer between domains.
@@
To enable positive while preventing negative transfer, we have pre-trained a transformer model on all data domains and then created fine-tuned models for individual domains. This approach improved predictive performance considerably and outperformed several existing methods, although the overall performance of all models still leaves room for improvement.
@@
Since modeling the binding of individual T-cell receptors (TCRs) and peptide:MHC complexes is difficult and does not directly capture the T-cell response (Hudson et al., 2023; Yin et al., 2022; Fischer et al., 2020), many prediction models, including ours, assume general patterns in peptide sequences that make peptide recognition more likely in many individuals (Wells et al., 2020; Ogishi and Yotsuyanagi, 2019). This assumption is based on the public TCR repertoire concept of antigen-recognizing TCRs being likely shared among individuals (Chen et al., 2017).
@@
Once more data of paired TCR and peptide-MHC complexes become available and protein structure predictions are further improved, the public TCR repertoire assumption might be relaxed. Modeling the interaction between individual TCRs and peptide-MHC complexes has the potential to improve T-cell response prediction in principle.
@@
In a broader context, the present study demonstrates the importance of knowing possibly hidden underlying structures in the training data for two reasons. First, the risk of shortcut learning is well known in domains like image recognition or natural language processing (Geirhos et al., 2020). However, in biological sequence data, invalid features might be harder to discover since they can be encoded in more complex patterns spanning only specific positions in sequences, as it is for anchor positions in peptides (see Figure 1b). Second, knowing about the heterogeneity of a data set and viewing it as a multi-domain structure can motivate the use of transfer learning techniques. However, as we have observed with ADA-T and as it is also reported for other tasks such as classification of images, texts, or molecules (Sagawa et al., 2022), transfer learning methods like adversarial domain adaptation can sometimes decrease the classification performance. Similarly, caution has to be applied in the data selection process since increasing a data set at the cost of more heterogeneity can also lead to instances of negative transfer.
@@
In a similar direction it is noteworthy that to this date, all approaches for T-cell response prediction use their own specific data selection criteria, differing, for example, in the selected assays, measured cytokines, or the considered peptide sources (Müller et al., 2023; Gfeller et al., 2023; Dhanda et al., 2018). We think that a benchmark data set that is well characterized in terms of its underlying structure, together with an evaluation scheme that is invariant to possible instances of shortcut learning, would be beneficial to facilitate objective development and a fair comparison of T-cell response prediction methods.

### v2

@@ 3.2 Negative transfer between domains
In a per-source setting, we train separate BASE-T models for every peptide source, each of which contains peptides from several MHC alleles. To make the performance of the models in the per-source setting directly comparable to the multi-domain setting, we first aggregate the predictions for the each source (performed by different models) into one data set again. Then we apply the evaluation on groups of peptide source/MHC allele combinations exactly as in the multi-domain setting. This evaluation also ensures that only predictions coming from the same per-source model are compared to each other because AUC values are computed per peptide source/MHC allele combination. Conversely, the per-allele setting is a combination of separate BASE-T models for each MHC allele, each of which is trained on data from several peptide sources.
The corresponding validation study (Supplement S5) shows that the AUC  of the multi-domain model is reduced by 0.05 compared to the per-source models and by 0.02 compared to the per-allele models, confirming our hypothesis that there is negative transfer between peptide sources and, to a smaller degree, between MHC alleles.
@@ 3.3 Improved predictions by per-source fine-tuning
FINE-T achieves higher AUC  values than BASE-T for all ten source/MHC combinations. We also observe that the difference is stronger for MHC I peptides than for MHC II peptides, which may be due to more room for improvement as BASE-T performs worse for MHC I compared to MHC II peptides. We also studied the performance of separate per-source models without fine-tuning. The resulting AUC values (see Supplement S6) are overall in between BASE-T and FINE-T, suggesting that the combination of both ideas, per source models and fine-tuning, are responsible for the superior performance of FINE-T.
By generating many random permutations of the T-cell response labels, we also verify that FINE-T learns a signal that cannot be explained by random structure in the data (see Supplement S7).
Figure 5: Test performance on human peptides presented on (a) MHC class I and (b) MHC class II. For both MHC classes, the first two rows correspond to existing models from the literature, which are trained on other data sets. Bag-of-AA baseline and FINE-T are trained on the same data. Blue bars correspond to the evaluation on the previously completely unused test set. Grey bars show the mean AUC from the final nested cross validation.
@@ 3.4 Comparison to other tools for human peptides
T-cell response predictions for peptides from human proteins are of special interest for the development of personalized cancer vaccines since neo-antigens also originate from human proteins [25]. Thus, we focus our analysis on human peptides and compare the performance of FINE-T, which was the best model in the validation study, to existing pre-trained models from the literature.
NetMHCpan 4.1 and NetMHCIIpan 4.0 [4] have been trained to predict the presentation of peptides on MHC molecules. Despite not being trained on T-cell response data, we include these models since they are commonly used in the context of personalized cancer vaccines. PRIME 2.0 [7] has been trained on T-cell response data with MHC I neo-antigens, whereas CD4Episcore [8] has been trained on MHC II T-cell response data from several peptide sources. Since the training data for all pre-trained models deviates from our training data, we also include the Bag-Of-AA baseline, trained on the same data set as FINE-T, in the comparison. In contrast to our previous studies, we now train a model on the union of the first four data subsets, and evaluate on the previously unused data subset $\mathcal{D}_{5}$. In addition, we also carry out a full nested cross validation. The results are shown in Figure 5.
For MHC I peptides, FINE-T achieves with a value of 0.615 the highest AUC among all models. PRIME 2.0 yields the lowest performance, which is likely due to the differences in the training data. For MHC II peptides, FINE-T also performs better than existing approaches, and the AUC is, with a value 0.68 even slightly better than in the validation study. While [26] have observed that combining MHC binding predictions from several alleles has a predictive value for immunogenicity, on the data set we use this applies only to MHC I (NetMHCpan 4.1) but not to MHC II (NetMHCIIpan 4.0). The large performance difference between the Bag-Of-AA model and FINE-T suggests that the possibility to use positional information of amino acids in the peptide sequence is indeed beneficial, and more so for MHC class II compared to MHC class I.
@@ 4 Discussion
In this work, we studied the effect of various transfer learning techniques on the problem of T-cell response prediction. We have observed that predicting T-cell responses to peptides is a challenging problem not only due to the limited sample size, but also since the data features a multi-domain structure, which entails, as illustrated in this article, the danger of shortcut learning.
@@
To ensure that performance estimates are not inflated by the use of shortcuts, we have defined a domain-aware evaluation scheme. We have then demonstrated that shortcut learning can be effectively reduced by applying adversarial domain adaptation [in an approach dubbed ADA-T. This shortcut reduction did, however, not directly improve predictive performance, which is likely due to large domain differences and negative transfer between domains.
@@
To enable positive while preventing negative transfer, we have pre-trained a transformer model on all data domains and then created fine-tuned models for individual domains, resulting in a model dubbed FINE-T. This approach improved predictive performance considerably and performs slightly better than several existing methods, although the overall performance of all models still leaves room for improvement.
We thus conclude that the overall effect on transfer learning as a broad strategy on T-cell response prediction is mixed, with FINE-T showing benefits, whereas ADA-T being not helpful. Beyond raw predictive performance, there a few other points to consider, though.
@@
Since modeling the binding of individual T-cell receptors (TCRs) and peptide:MHC complexes is difficult and does not directly capture the T-cell response [27, 28, 29], many prediction models, including ours, assume general patterns in peptide sequences that make peptide recognition more likely in many individuals [30, 11]. This assumption is based on the public TCR repertoire concept of antigen-recognizing TCRs being likely shared among individuals [31].
@@
Once more data of paired TCR and peptide-MHC complexes become available and protein structure predictions are further improved, the public TCR repertoire assumption might be relaxed. Modeling the interaction between individual TCRs and peptide-MHC complexes has the potential to improve T-cell response prediction in principle [32].
@@
In a broader context, the present study demonstrates the importance of knowing possibly hidden underlying structures in the training data for two reasons. First, the risk of shortcut learning is well known in domains like image recognition or natural language processing [18]. However, in biological sequence data, invalid features might be harder to discover since they can be encoded in more complex patterns spanning only specific positions in sequences, as it is for anchor positions in peptides (see Figure 1b). Second, knowing about the heterogeneity of a data set and viewing it as a multi-domain structure can motivate the use of transfer learning techniques. However, as we have observed with ADA-T and as it is also reported for other tasks such as classification of images, texts, or molecules [33], transfer learning methods like adversarial domain adaptation can sometimes decrease the classification performance. Similarly, caution has to be applied in the data selection process since increasing a data set at the cost of more heterogeneity can lead to instances of negative transfer [12].
@@
In a similar direction it is noteworthy that to this date, all approaches for T-cell response prediction use their own specific data selection criteria, differing, for example, in the selected assays, measured cytokines, or the considered peptide sources [34, 7, 8]. We think that a benchmark data set that is well characterized in terms of its underlying structure, together with an evaluation scheme that is invariant to possible instances of shortcut learning, would be beneficial to facilitate objective development and a fair comparison of T-cell response prediction methods.
[…]
Supplementary information
There is one file with supplementary material attached to the submission.
Declarations
@@
  • Conflict of interest
    No competing interest is declared.
  • Ethics approval and consent to participate
    Not applicable
  • Consent for publication
    Not applicable
  • Data availability
    Data is available under github.com/JosuaStadelmaier/T-cell-response-prediction.
  • Materials availability
    Not applicable
  • Code availability
    Source code and trained models are available under github.com/JosuaStadelmaier/T-cell-response-prediction.
  • Author contribution
    J.S., B.M., and R.E conceived the experiment(s), J.S. implemented the software, J.S. conducted the experiment(s), J.S., B.M., and R.E. analysed the results. J.S., B.M. and R.E. wrote and reviewed the manuscript.